# Explainable AI for Curie Temperature Prediction in Magnetic Materials

### v1

@@ Abstract
We explore machine learning techniques for predicting Curie temperatures of magnetic materials using the NEMAD database. By augmenting the dataset with composition-based and domain-aware descriptors, we evaluate the performance of several machine learning models. We find that the Extra Trees Regressor delivers the best performance reaching an R2 score of up to 0.85 $\pm$ 0.01 (cross-validated) for a balanced dataset. We employ the k-means clustering algorithm to gain insights into the performance of chemically distinct material groups. Furthermore, we perform the SHAP analysis to identify key physicochemical drivers of Curie behavior, such as average atomic number and magnetic moment. By employing explainable AI techniques, this analysis offers insights into the model’s predictive behavior, thereby advancing scientific interpretability.
@@ 1  Introduction
Curie temperature (TC) is a critical property that defines the transition of a material from a ferromagnetic to a paramagnetic state. Predicting TC with high accuracy is essential for the discovery of magnetic materials for data storage, spintronics, and energy applications. Traditional approaches based on quantum mechanical computations or empirical models are often limited in scalability and accuracy. In recent years, machine learning (ML) has emerged as a promising alternative for property prediction across materials science domains [1, 2, 3, 4, 5].
Building on this momentum, several works have proposed the use of ML models trained on curated magnetic datasets. In particular, the recent study [6] introduced the NEMAD database, which aggregates experimentally measured magnetic transition temperatures and compositions. Similarly, the study by Belot et al. [7] utilized two of the largest available datasets of experimental Curie temperatures—comprising over 2,500 materials for training and more than 3,000 entries for validation—to compare machine learning strategies for predicting Curie temperature solely from chemical composition.
@@ 2  Dataset Preparation and Model Evaluation
We begin our study with the NEMAD database [6], a publicly available dataset compiled using large language models, comprising 26,706 magnetic materials with records for chemical composition, phase transition temperatures (e.g., Curie or Néel temperatures), structural characteristics, and related magnetic properties. We pre-processed the database to remove ambiguous entries, duplicates, and entries with missing or non-numeric temperatures. We also standardized chemical formulas using the pymatgen library.
@@
The Extra Trees algorithm achieved the best overall performance, with an $R^{2}$ value of 0.87, MAE of 54 K, and RMSE of 106 K. This performance surpasses that of the Random Forest algorithm and other models considered. With a balanced dataset we were able to achieve $R^{2}\sim 0.85$ with a standard deviation of $0.01$. The balanced dataset was obtained by randomly removing around 4% of samples with Curie temperatures less than 300 K [6].
The Extra Trees (Extremely Randomized Trees) algorithm employs an ensemble learning method similar to the Random Forests model, but with two notable distinctions:
@@
(a) SHAP bar plot showing the average impact of each feature on model output.
(b) SHAP beeswarm plot showing feature influence per sample. Color indicates feature value.
Figure 3: SHAP analysis of the ExtraTrees model. (a) shows the average contribution of each feature across the dataset, while (b) reveals the variation of feature impacts for individual predictions.
3  SHAP Analysis
To interpret the predictions of our Extra Trees model in the context of Explainable AI, we use SHAP (SHapley Additive exPlanations). SHAP assigns each feature an importance value for a particular prediction, based on cooperative game theory. This allows us to quantify how much each input feature contributes to increasing or decreasing the predicted Curie temperature.
SHAP Bar Plot: Figure (3), left panel, shows the SHAP summary bar plot, which ranks the features by their average absolute SHAP values. This plot reveals that the Mean Magnetic Moment is the most influential feature. This descriptor captures the average magnetic moment of constituent atoms, which is strongly linked to Curie temperature, due to its role in long-range magnetic ordering. Features like Average Deviation of Number of Unfilled f-Orbitals and Mean Ground State Volume/Atom also show significant contributions, reflecting magnetic configuration and spatial arrangement of electrons.
Interestingly, elemental fractions like Fraction of Fe and Fraction of Mn have moderate importance, suggesting that while the presence of individual elements is informative, more complex electronic or structural descriptors may serve as stronger predictors of Curie temperature.
SHAP Beeswarm Plot: The SHAP beeswarm plot Figure (3), right panel, illustrates the distribution of SHAP values for each feature across all samples. Each point represents a single material, with color indicating the corresponding feature value. This visualization reveals not only which features are most influential, but also how they affect the model’s predictions. For example, higher values of the average magnetic moment of constituent atoms tend to increase the predicted Curie temperature, while lower values generally lead to decreased predictions.
@@ 3  SHAP Analysis
We can see from the plot that the Mean Magnetic Moment, which represents the average magnetic moment of the ground-state structure, exhibits a strong positive influence when its value is high (shown in red), shifting SHAP values to the right. This suggests that materials with higher magnetic moments are strongly associated with higher Curie temperatures, which is physically intuitive.
@@
In contrast, the average deviation of magnetic moment shows a somewhat different pattern. High values (in red) are generally associated with negative SHAP values, indicating a suppressive effect on Curie temperature. This indicates that when the distribution of atomic magnetic moments within a compound is more uneven, the model predicts a lower Curie temperature.
Additionally, the fraction of Mn feature, representing the proportion of manganese in the compound, shows a symmetrical SHAP value distribution with a subtle trend: higher Mn fractions (red) tend to push SHAP values upward. This behavior reinforces manganese’s well-established role in stabilizing magnetic ordering, which aligns with its frequent presence in high-$T_{C}$ magnetic materials.
These patterns not only validate domain expectations but also reveal hidden non-linearities that traditional feature importance metrics might overlook.
Figure 4: Elbow plot showing the inertia values for different numbers of clusters. The curve begins to flatten around $k=4$, suggesting a reasonable balance between variance capture and model complexity.
@@ 4  K-means Clustering Analysis
Figure 5: Figure 3(a) visualizes the dominant elements in each cluster, reinforcing chemically meaningful groupings. Figure 3(b) shows the t-SNE visualization for the dataset, where each point represents a material colored by its assigned cluster. The separation among clusters, particularly the distinct distribution of Cluster 1, suggests notable differences in structural or compositional properties that impact Curie temperature prediction.
@@ 6  Acknowledgments
The code and data for this project is available on github: https://github.com/dradeelajaib/Curie-Prediction

### v3

@@ Abstract
We explore machine learning techniques for predicting Curie temperatures of magnetic materials using the NEMAD database. By augmenting the dataset with composition-based and domain-aware descriptors, we evaluate the performance of several machine learning models. We find that the Extra Trees Regressor delivers the best performance, reaching an R2 score of up to 0.85 $\pm$ 0.01 (cross-validated) for a balanced dataset. We employ the k-means clustering algorithm to gain insights into the performance of chemically distinct material groups. Furthermore, we perform SHAP analysis to identify key physicochemical drivers of Curie behavior, such as average atomic number and magnetic moment. By employing explainable AI techniques, this analysis offers insights into the model’s predictive behavior, thereby advancing scientific interpretability.
@@ 1  Introduction
Curie temperature ($T_{c}$) is a critical property that defines the transition of a material from a ferromagnetic to a paramagnetic state. Predicting $T_{c}$ with high accuracy is essential for the discovery of magnetic materials for data storage, spintronics, and energy applications. Traditional approaches based on quantum mechanical computations or empirical models are often limited in scalability and accuracy. In recent years, machine learning (ML) has emerged as a promising alternative for property prediction across materials science domains [7, 2, 3, 1, 9, 4, 8, 5, 6].
Building on this momentum, several recent studies have proposed the use of ML models trained on curated magnetic datasets. In particular, the recent study [10] introduced the NEMAD database, which aggregates experimentally measured magnetic transition temperatures and compositions. Similarly, the study by  [11] utilized two of the largest available datasets of experimental Curie temperatures—comprising over 2,500 materials for training and more than 3,000 entries for validation—to compare machine learning strategies for predicting Curie temperature solely from chemical composition.
@@ 2  Dataset Preparation and Model Evaluation
We begin our study with the NEMAD database [10], a publicly available dataset compiled using large language models, comprising 26,706 magnetic materials with records for chemical composition, phase transition temperatures (e.g., Curie or Néel temperatures) and related magnetic properties. We preprocessed the database to remove ambiguous entries, duplicates, and entries with missing or non-numeric temperatures. We also standardized chemical formulas using the pymatgen library.
From the original 33,668 entries in the NEMAD dataset, we retained 21,024 samples (62.4%). Rows with missing material names or Curie temperatures (36.6%) were removed. Curie temperatures were parsed to extract numerical Kelvin values, and entries with non-parsable or non-physical temperatures (approximately 1.5% of the data) were discarded. This generated a clean dataset for the machine learning analysis.
@@
The Extra Trees algorithm achieved the best overall performance, with an $R^{2}$ value of 0.87, MAE of 54 K, and RMSE of 106 K. This performance surpasses that of the Random Forest algorithm and other models considered. With a balanced dataset, we were able to achieve $R^{2}\sim 0.85$ with a standard deviation of $0.01$. The balanced dataset was obtained by randomly removing around 4% of samples with Curie temperatures less than 300 K [10].
The Extra Trees (Extremely Randomized Trees) algorithm employs an ensemble learning method similar to the Random Forest model, but with two notable distinctions:
@@
(a) Cluster-wise elemental composition percentages.
(b) t-SNE projection with cluster labels.
Figure 3: SHAP analysis of the Extra Trees model. Panel (a) shows SHAP bar plot showing the mean absolute SHAP value of the top features contributing to Curie temperature predictions across the entire dataset. Higher bars indicate stronger global influence on model output. Panel (b) displays the SHAP beeswarm plot illustrating the distribution of feature impacts for individual samples. Each point represents a material. Feature strength is encoded in color (red = high, blue = low), showing whether large or small values increase or decrease the predicted Curie temperature.
2.1  SHAP Analysis
To better understand the model’s behavior and identify the most influential features contributing to the prediction of Curie temperature, we performed SHAP (SHapley Additive exPlanations) analysis. SHAP values offer a model-agnostic, consistent way to explain predictions by attributing importance scores to individual features.
SHAP Bar Plot: Figure 3 shows the SHAP summary bar plot of the top 10 features based on their mean absolute SHAP values. Among the dominant predictors, the mean magnetic moment stands out, reinforcing the well-established role of atomic-scale magnetism in determining $T_{c}$. The fraction of Fe also shows high importance, which aligns with its prevalence in high-$T_{c}$ compounds and well-known ferromagnets. Features such as the mode magnetic moment, max magnetic moment, and fraction of Co all rank highly, further confirming that the magnetic nature of the constituent atoms is a primary driver of the model’s predictions.
Interestingly, features such as rare earth proportion, mean ground state volume per atom, and group number deviation appear in the top contributors. These descriptors capture the complexity of atomic arrangements and compositional heterogeneity, both of which are known to modulate magnetic interactions in complex systems.
@@ 2.1  SHAP Analysis
SHAP Beeswarm Plot: The SHAP beeswarm plot in Figure 3, right panel, illustrates the distribution of SHAP values for each feature across all samples. Each point represents a single material, with color indicating the corresponding feature value. This visualization reveals not only which features are most influential, but also how they affect the model’s predictions. For example, we can see from the plot that the Mean Magnetic Moment, which represents the average magnetic moment of the ground-state structure, exhibits a strong positive influence when its value is high (shown in red), shifting SHAP values to the right. This suggests that materials with higher magnetic moments are strongly associated with higher Curie temperatures, which is physically intuitive.
@@
In contrast, the f-Orbital Deviation shows a somewhat different pattern. High values (in red) are generally associated with negative SHAP values, indicating a suppressive effect on Curie temperature. This deviation metric captures the compositional spread in elements possessing f-electrons, such as rare-earth or actinide elements.
Taken together, these plots indicate that the model captures meaningful physical trends: magnetic moment-related features dominate, while elemental diversity exhibits subtle patterns. This supports the model’s reliability and its alignment with known physics.
Figure 4: Elbow plot showing the inertia values for K-means clustering as a function of the number of clusters $k$. A rapid decrease in inertia is observed from $k=1$ to $k=3$, followed by a clear flattening beyond $k=4$, indicating that adding more clusters yields only marginal improvement in explaining the data structure. This behavior supports the choice of $k=4$ as a balanced and physically meaningful clustering configuration for the Curie temperature dataset.
@@ 3  K-means Clustering Analysis
Figure 5: Figure 5: (a) Cluster-wise elemental composition percentages highlighting the dominant chemical elements within each K-means cluster. The distributions reveal clear chemical themes, such as Fe/Co-rich metallic systems, oxide-based compositions enriched in O, Mn, and La, and Mn/Ni-containing intermetallics. (b) t-SNE projection of the high-dimensional descriptor space into two dimensions, with each point representing a material colored by its assigned cluster. The distinct separation—especially the well-isolated oxide-rich Cluster 1—demonstrates that the clustering captures meaningful compositional differences relevant to Curie temperature variation.
@@
Cluster 0
Cluster 1
Cluster 2
Cluster 3
Figure 6: SHAP beeswarm plots for the four K-means clusters, showing how feature values contribute to Curie temperature predictions within each chemical subgroup. Cluster 0 (top left) contains Fe/Co-rich materials where magnetic-moment descriptors are particularly strong contributors. Cluster 1 (top right) includes oxide-based compounds, where magnetic-moment features remain important but orbital-deviation and magnetic-variance descriptors gain greater influence. Cluster 2 (bottom left), composed of Mn/Ni intermetallics, is primarily driven by magnetic-moment and atomic-size descriptors. Cluster 3 (bottom right) is a smaller, compositionally diverse group with a relatively broader SHAP contribution.
4  Cluster-wise SHAP Analysis
To further enhance the interpretability of our machine learning models for Curie temperature ($T_{c}$), we applied the SHAP analysis for each of the four clusters identified via KMeans. By isolating materials into chemically similar groups, we can uncover subtle relationships and also better understand which descriptors govern $T_{c}$ in different regimes.
Cluster 0, which is the largest cluster (4,744 materials), contains a significant number of Fe-, Co-, and rare-earth-based compounds. The model achieves strong performance with MAE = 86.83 K and $R^{2}=0.80$. As shown in top left panel of Figure 6, the most influential features include the mean magnetic moment, mode magnetic moment, and fraction of Co, all positively correlated with $T_{c}$. These results reflect known physical principles — high magnetic moments enhance exchange interactions and stabilize ferromagnetic ordering. Additional features such as minimum Mendeleev number and mode column suggest that periodic trends play a secondary yet non-trivial role.
Cluster 1 (4,826 materials) includes many oxide-based perovskites rich in Mn, La, and O. The model performance is lower (MAE = 100.21 K, $R^{2}=0.58$), likely due to the structural and chemical complexity of these materials. Figure 6 (top right plot) reveals that fraction of Fe, average deviation of magnetic moment, and f-orbital deviation are key predictive features. The prominence of orbital and magnetic variance descriptors suggests the presence of competing magnetic mechanisms such as double exchange and Jahn-Teller distortions. These subtleties may explain the slightly reduced predictive power compared to Cluster 0.
Cluster 2 (2,730 materials) consists of Mn- and Ni-based intermetallics, including Heusler-like alloys. Despite compositional richness, the model achieves good performance (MAE = 88.68 K, $R^{2}=0.73$). In Figure 6 (bottom left plot), features such as mean magnetic moment, fraction of Mn, and maximum covalent radius dominate. This aligns with the tunability of intermetallic compounds, where structural and electronic changes can significantly influence magnetic ordering. The importance of spatial descriptors further highlights the role of atomic packing and size effects in modulating $T_{c}$.
Cluster 3 is the smallest cluster (760 materials) and exhibits the poorest performance (MAE = 138.88 K, $R^{2}=0.38$). This may be due to its small size and its unique chemical composition. As seen in Figure 6 (bottom right plot), while magnetic descriptors still lead, there is greater importance given to features such as La fraction, Sr fraction, and f-orbital deviation. These may hint at more subtle magnetic phenomena that our current descriptor set cannot fully capture. The high variance in SHAP values also suggests that no single feature consistently governs $T_{c}$ across the cluster, pointing to the need for more tailored descriptors or domain-specific models.
SHAP analysis across clusters confirms that magnetic moment descriptors are universally important predictors of Curie temperature. However, the prominence of different supporting features (e.g., orbital variances, group numbers, atomic radii) shifts based on the chemical and structural nature of the materials within each cluster. This highlights that even with a shared target, distinct materials demand distinct explanatory regimes.
We conclude that cluster-specific modeling not only improves performance but also enriches physical insight. Future work may benefit from hybrid clustering methods (e.g., hierarchical or domain-guided) and the integration of additional physics-based descriptors tailored to each family. It is also worth noting that while SHAP helps explain model behavior, it doesn’t always align with physical intuition, especially for complex oxides. This is expected, as SHAP reflects correlations based on available descriptors, which may miss deeper physics. Its insights are valuable, but approximate.
@@ 6  Acknowledgments
The code and data for this project are available on GitHub: https://github.com/dradeelajaib/Curie-Prediction